OpenStudy (jacobbirddog):

Please Help! Will Fan And Medal! 5 = –(–z + 3)

OpenStudy (anonymous):

-3 times -3 minus 4

OpenStudy (anonymous):

5=−(−z+3) I will try to solve your equation. 5=z−3(Simplify both sides of the equation) 5+−1z=z−3+−1z(Add -1z to both sides) −z+5=−3 −z+5+−5=−3+−5(Add -5 to both sides) −z=−8 (−z) (−1) = (−8) (−1) (Divide both sides by -1) z=8